# Full Event Particle-Level Unfolding with Variable-Length Latent Variational Diffusion

### v1

@@ 4.2 Standard Model Results
The VLD model performance is evaluated on a testing sample, generated identically to the training sample. Figures 5 and  6 show the kinematic distributions of the leptons and jets respectively [^2] Excepting the lepton masses, whose values are well known and need not be unfolded.. The distributions are inclusive over all leptons or jets in the events. Distributions labelled “truth” are from events without detector effects, the target of the unfolding. Distributions labeled “detector” are from events which include simulated effects of the detector and are used to condition generation. The distributions labelled “unfolded” are from particle level events generated by the VLD unfolding algorithm.
@@
In general, there is excellent agreement between the unfolded and truth distributions for the inclusive lepton and jet kinematics. This is expected, as the network is trained to directly predict these quantities. Some disagreement is found at the kinematic edges, for example at low $p_{\mathrm{T}}$ and mass or at extreme values of $\eta$, due to a lack of examples of events migrating across the selection requirements from particle to detector level. This could be mitigated by imposing a tighter selection in generation than in training, avoiding the need to learn sharp cut-offs in the target distributions. Table 1 displays three measures of distance [^3] Measure definitions are in Appendix C of Ref. [18]. between the truth distributions and the corresponding detector and unfolded distributions for the jet and lepton kinematics. In most cases, the distance to the truth distribution is smaller for the unfolded distributions than the detector distributions. One exception is in the jet $\eta$ observable, due to events migrating across the event selection between particle and detector levels at high $|\eta|$.
@@
Figure 10 shows the kinematic distributions for the reconstructed hadronically-decaying top quark, leptonically-decaying top quark, and the $t\bar{t}$ system. Closure in these distributions is not as good as for the kinematics of the jets and leptons, though it is not surprising that these observables are more difficult to model as they were not directly used to optimize the networks. In particular, the top quark mass distributions are very difficult to unfold, with the sharp peaks at truth level not reproduced after unfolding. This is an illustration of the well known difficulty in modeling sharp features with generative models [12].
@@ 4.3 Performance on Dataset with BSM Physics Injection
Figure 13 shows the truth, detector, and unfolded lepton and jet kinematics and multiplicity in the SM and EFT samples. There is good agreement almost everywhere between the unfolded and truth distributions for the EFT sample despite large differences with the SM training sample. The reconstructed top quark and $t\bar{t}$ system kinematics are shown in Fig. 14. They show similar disagreements in top $p_{\mathrm{T}}$ and mass as was observed in unfolding the SM sample, which are much smaller than the differences between the SM and EFT truth distributions. This indicates that the choice of prior training sample has not induced a significant bias in the unfolding derived using VLD, at least in this test case. Nonetheless, it may be necessary to apply the iterative method proposed in [14] to remove all dependence on the prior when applying to real data. A full set of comparisons between the EFT and SM samples can be found in Appendix B.
@@ 5 Outlook and Discussion
VLD was able to accurately predict truth-level distributions in the vast majority of phase space. However, mis-modelling was observed in regions where the truth-level distributions had sharp features or limited statistics in the training data. Previous work [18] accounted for the sharply peaked partonic top-quark mass distributions using a physics-inspired consistency loss term that ensured the predicted 4-vector components maintained the correct correlations. Here, this loss was not necessary to ensure consistency in the predicted 4-vectors for the particle-level leptons and jets. Mis-modelling was however observed in the sharply peaked particle-level top quark mass distributions. Since the particle-level top quark masses are reconstructed post-unfolding and VLD is not directly optimized to generate these distributions, applying a similar physics consistency loss is non-trivial. Other physics constraints, such as those enforcing symmetries or conversation laws, could also be incorporated into the objective in principle. Such physics constraints can easily be added to the loss function used to optimize the VAE; however adding such loss terms to the diffusion process would require performing computationally expensive inference at training time. Table 5 shows the distance metrics between detector and unfolded distributions for a selection of observables, distinguishing between the contributions to post-unfolding distance from the VAE or the diffusion parts of the network. The majority of the error comes from the diffusion process, and so an investigation of how best to utilize physics constraints is left to future work. Full error breakdown tables can be found in Appendix C.
@@ 6 Conclusions
In general, there is excellent closure between truth and unfolded distributions, both for a sample similar to the training sample, and an alternative. This lack of prior dependence strongly motivates the use of VLD for unfolding.
Some mis-modeling of kinematic distributions is seen near edges of the selection, due to a lack of training samples. Potential mitigation strategies were discussed in Section 5. Observables not included in the VLD training, such as the reconstructed top quark kinematics, showed larger mis-modeling. Most significant is the mis-modeling of the top quark mass distributions, which are sharply peaked at particle-level and not well reconstructed at detector level. Future work may incorporate additional physics constraints to better handle these kinds of observables.
@@ Appendix A Variable Definitions
Appendix B Full set of EFT Distributions
@@
Appendix C Full Error Breakdown Tables

### v2

@@ 4.2 Standard Model Results
The VLD model performance is evaluated on a testing sample, generated identically to the training sample. Figures 5 and  6 show the kinematic distributions of the leptons and jets respectively [^2] Excepting the lepton masses, whose values are well known and need not be unfolded.. The distributions are inclusive over all leptons or jets in the events. Distributions labelled “truth” are from events without detector effects, the target of the unfolding. Distributions labeled “detector” are from events which include simulated effects of the detector and are used to condition generation. The distributions labelled “unfolded” are from particle level events generated by the VLD unfolding algorithm, and the shaded error bands on the unfolded distributions are obtained by sampling each event 128 times. The ability to sample the generative model multiple times for a single detector-level event is an additional benefit of the generative approach [69]. However in this application, the uncertainties obtained from sampling the model are strictly larger than the statistical uncertainties in the distributions.
@@
In general, there is excellent agreement between the unfolded and truth distributions for the inclusive lepton and jet kinematics. This is expected, as the network is trained to directly predict these quantities. Some disagreement is found at the kinematic edges, for example at low $p_{\mathrm{T}}$ and mass or at extreme values of $\eta$, due to a lack of examples of events migrating across the selection requirements from particle to detector level. This could be mitigated by imposing a tighter selection in generation than in training, avoiding the need to learn sharp cut-offs in the target distributions. Table 1 displays three measures of distance between the truth distributions and the corresponding detector and unfolded distributions for the jet and lepton kinematics. The definitions of these measures are presented in Appendix B. In most cases, the distance to the truth distribution is smaller for the unfolded distributions than the detector distributions. One exception is in the jet $\eta$ observable, due to events migrating across the event selection between particle and detector levels at high $|\eta|$.
Table 1: Wasserstein, Energy, and Kulback-Leibler distance measures between truth and detector or truth and unfolded distributions for the jet kinematics (top), lepton kinematics (middle), and event-level observables (bottom). The unfolded distribution uncertainties are estimated by sampling each event 128 times.
@@
Figure 10 shows the kinematic distributions for the reconstructed hadronically-decaying top quark, leptonically-decaying top quark, and the $t\bar{t}$ system. Closure in these distributions is not as good as for the kinematics of the jets and leptons, though it is not surprising that these observables are more difficult to model as they were not directly used to optimize the networks. In particular, the top quark mass distributions are very difficult to unfold, with the sharp peaks at truth level not reproduced after unfolding. This is an illustration of the well known difficulty in modeling sharp features with generative models [12]. Corner plots illustrating correlations between these observables in the truth and predicted distributions are provided in Appendix E. These show that correlations between arbitrary observables are faithfully captured by VLD, despite the difficulty in modeling some of the marginal distributions.
@@ 4.3 Performance on Dataset with BSM Physics Injection
Figure 13 shows the truth, detector, and unfolded lepton and jet kinematics and multiplicity in the SM and EFT samples. There is good agreement almost everywhere between the unfolded and truth distributions for the EFT sample despite large differences with the SM training sample. The reconstructed top quark and $t\bar{t}$ system kinematics are shown in Fig. 14. They show similar disagreements in top $p_{\mathrm{T}}$ and mass as was observed in unfolding the SM sample, which are much smaller than the differences between the SM and EFT truth distributions. This indicates that the choice of prior training sample has not induced a significant bias in the unfolding derived using VLD, at least in this test case. Nonetheless, it may be necessary to apply the iterative method proposed in [14] to remove all dependence on the prior when applying to real data. A full set of comparisons between the EFT and SM samples can be found in Appendix C.
@@ 5 Outlook and Discussion
VLD was able to accurately predict truth-level distributions in the vast majority of phase space. However, mis-modelling was observed in regions where the truth-level distributions had sharp features or limited statistics in the training data. Previous work [18] accounted for the sharply peaked partonic top-quark mass distributions using a physics-inspired consistency loss term that ensured the predicted 4-vector components maintained the correct correlations. Here, this loss was not necessary to ensure consistency in the predicted 4-vectors for the particle-level leptons and jets. Mis-modelling was however observed in the sharply peaked particle-level top quark mass distributions. Since the particle-level top quark masses are reconstructed post-unfolding and VLD is not directly optimized to generate these distributions, applying a similar physics consistency loss is non-trivial. Other physics constraints, such as those enforcing symmetries or conversation laws, could also be incorporated into the objective in principle. Such physics constraints can easily be added to the loss function used to optimize the VAE; however adding such loss terms to the diffusion process would require performing computationally expensive inference at training time. Table 5 shows the distance metrics between detector and unfolded distributions for a selection of observables, distinguishing between the contributions to post-unfolding distance from the VAE or the diffusion parts of the network. The majority of the error comes from the diffusion process, and so an investigation of how best to utilize physics constraints is left to future work. Full error breakdown tables can be found in Appendix D.
@@ 6 Conclusions
In general, there is excellent closure between truth and unfolded distributions, both for a sample similar to the training sample, and an alternative. Some mis-modeling of kinematic distributions is seen near edges of the selection, due to a lack of training samples. Potential mitigation strategies were discussed in Section 5. Observables not included in the VLD training, such as the reconstructed top quark kinematics, showed larger mis-modeling. Most significant is the mis-modeling of the top quark mass distributions, which are sharply peaked at particle-level and not well reconstructed at detector level. Future work may incorporate additional physics constraints to better handle these kinds of observables.
@@
To encourage further development in generative unfolding methods, the datasets used are made publicly available [76]. The code base is available at https://github.com/Alexanders101/LVD/tree/main.
@@ Appendix A Variable Definitions
Appendix B Distance Metrics
As the parton global distributions do not have a known family of distributions to describe their components, model-free measures of distribution distance must be used to evaluate the models. Three different families of distance measures are used. These non-parametric distances are only defined for 1-dimensional distributions. As there is no commonly accepted way of measuring distance for $N$-dimensional distributions, the 1-dimensional distances are simply summed across the components. Although not ideal, it is enough to compare different models and rank them based on performance.
B.1 Wasserstein Distance
The Wasserstein distance, often referred to as the earth-mover distance, quantifies the amount of work it takes to reshape one distribution into another. This concept originated from the field of optimal transport and has found wide applications in many areas, including machine learning. An equivalent definition defines this distance as the minimum cost to move and transform the mass of one distribution to match another distribution. For a pair of 1-dimensional distribution samples, denoted $u$ and $v$, the Wasserstein distance can be computed in a bin-independent manner. This is achieved by computing the integral of the absolute difference between their empirical cumulative distribution functions (CDFs), $U(x)$ and $V(x)$.
[…]
B.2 Energy Distance
Energy distance is another statistical measure used to quantify the difference between two probability distributions based on empericial CDFs. It compares the expected distance between random variables drawn from the same distribution (intra-distribution) with the expected distance between random variables drawn from different distributions (inter-distribution). The Energy distance may be defined as the squared variant of the Wasserstein distance.
[…]
B.3 Kolmogorov-Smirnov Test
The two-sample Kolmogorov-Smirnov (K-S) test is a non-parametric statistical hypothesis test used to compare the underlying probability distributions of two independent samples. It is particularly useful in machine learning applications where the goal is to assess whether two datasets come from the same distribution or if they differ significantly, without making any assumptions about the underlying distribution shape. It is also based on empirical CDFs.
B.4 KL-Divergence
An alternative approach to empirical CDF approaches is to bin the data into histograms and compute discrete distribution distances from these histograms. The common Kullback–Leibler distance is used with three different bin sizes. After finding the histograms with $N$ bins for $1\leq i\leq N$, $P_{N}(i)$ and $Q_{N}(i)$, the discrete KL divergence is computed as
$D_{KL,N}=\sum_{i=1}^{N}P_{N}(i)\log{\left(\frac{P_{N}(i)}{Q_{N}(i)}\right)}$
Appendix C Full set of EFT Distributions
@@
Appendix D Full Error Breakdown Tables
@@
Appendix E Corner Plots of Top-Quark Kinematics
Figure 21: Corner plot illustrating the correlations between six dimensions that characterise the predicted hadronic top quark kinematics for the SM dataset.
Figure 22: Corner plot illustrating the correlations between six dimensions that characterise the truth hadronic top quark kinematics for the SM dataset.
Figure 23: Corner plot illustrating the correlations between six dimensions that characterise the predicted leptonic top quark kinematics for the SM dataset.
Figure 24: Corner plot illustrating the correlations between six dimensions that characterise the truth leptonic top quark kinematics for the SM dataset.